How do I find the value of x for this equation below
x+2x+5=x+19
3x-x=19-5
2x=14
x=7

Found 2 solutions by checkley71, razz288:
Answer by checkley71(8403)

x+2x+5=x+19
3x-x=19-5
2x=14
x=14/2
x=7
proof
7+2*7+5=7+19
7+14+5=26
26=26

Answer by razz288(1)

x+2x+5=x+19
3x-x=19-5
2x=14
x=7